# ViIK: Flow-based Vision Inverse Kinematics Solver with Fusing Collision Checking

### v1

@@ III-A Target Distribution
$\left\{\begin{array}[]{ll}p_{C_{c-f}}(c|p)&\\[2.0pt] p_{X^{\prime}}(x^{\prime}|p)\end{array}\right.,$ (5)
where $x^{\prime}=(c,x_{c})$, $x_{c}$ is whether collision-with-env occurs. The first one $p_{C_{c-f}}(c|p)$ is to train models without negative samples where models might perform badly near the boundary. In this paper, we choose $p_{X}((c,x_{c})|p)$ as the target distribution. After adding self-collision in the same way, the target distribution is:
@@ IV Experiments
IV-A Runtime Compared with the Classical Method
We compare the runtime of ViIK to the classical method, TRAC-IK+CC. The tolerance of TRAC-IK is set to $1\times 10^{-3}$ which is similar to the errors of ViIK shown in Tab. I. Moreover, Mesh Collider is used for collision checking in TRAC+CC. We let TRAC-IK repeatedly execute until the number of output solutions is the same as ViIK, $t_{ik}$ denotes the total runtime of TRAC-IK. Then we run collision checking for every configuration output by TRAC-IK, $t_{cc}$ denotes the total runtime of collision checking. The runtime of TRAC-IK+CC is:
@@ IV-A Runtime Compared with the Classical Method
TABLE I: Position and Angular Errors on the test set (150,000 random IK problems of Panda manipulator).
@@
IV-B Position Error and Angular Error Compared with the Learning-based Methods
We evaluate ViIK on the test set with 150,000 random IK problems of the Panda manipulator against IKFlow and PaddingFlow. We let ViIK, IKFlow, and PaddingFlow output 1000 solutions for each target pose. In Tab. I, three models (ViIK-2, ViIK-5, and ViIK-10 trained on the data of 2, 5, and 10 scenes respectively) can output competitive solutions compared with existing learning-based methods. ViIK outperforms IKFlow and PaddingFlow on position error which is about 1/2 of IKFlow, but on angular error, ViIK performs worse than PaddingFlow.
TABLE II: Self-collision and Collision-with-Env Rates on the test set (10,000 random IK problems of Panda manipulator).
@@ IV-B Position Error and Angular Error Compared with the Learning-based Methods
IV-C Self-collision and Collision-with-Env Rates
We present ViIK rates of self-collision and collision with the environments on the test sets. We test self-collision and collision-with-env rates on the test set with 10,000 random IK problems of the Panda manipulator. For each target pose, ViIK outputs 1000 solutions. For different scenes, we evaluate ViIK on the same test set but with different images of each scene. We report self-collision and collision-with-env rates over 10,000 target poses, as shown in Tab. II. The self-collision rates of ViIK are lower than 2%, which is significantly lower than the test set. As for the collision-with-env rates, ViIK can perform well in most scenes but is not ideal in some hard scenes (Env-1, and Env-5).
@@ V-B Experimental details
All models are GLOW-based[18] with 24 coupling layers. The training set has 5 million random samples in the C-space with the corresponding end-effector poses and 4882 sets of multi-view images of each environment. The batch size is 1024. The Adamw optimizer is used for training all models. The hyperparameters for training ViIK are shown in Tab. III. Moreover, the multi-views used in this paper are presented in Fig. 6 with an example in Env-1.
TABLE III: Hyperparameters for training ViIK.

### v2

@@ III-A Target Distribution
$\left\{\begin{array}[]{ll}p_{1}=p_{C_{c-f}}(c|p)&\\[2.0pt] p_{2}=p_{C}(c|p,x_{c})&\\[2.0pt] p_{3}=p_{X^{\prime}}(x^{\prime}|p)\end{array}\right.,$ (5)
where $x^{\prime}=(c,x_{c})$, $x_{c}$ is whether collision-with-env occurs. The first one $p_{C_{c-f}}(c|p)$ is to train models without negative samples where models might perform badly near the boundary. As for the second one $p_{C}(c|p,x_{c})$, the model has similar performance on position and angular error compared with $p_{X}((c,x_{c})|p)$, but performs badly on collision-sith-env and self-collision rates, as shown in Tab. I. In this paper, we use $p_{X}((c,x_{c})|p)$ as the target distribution. After adding self-collision in the same way, the target distribution is:
@@ III-B2 Loss Function
TABLE I: Target distributions comparison on the test set (150,000 random IK problems of Panda manipulator).
Model Position Error (mm, $\downarrow$) Angular Error (deg, $\downarrow$) Collision-with-Env Rates (%, $\uparrow$) Self-collision Rates (%, $\uparrow$)
Env-1 $p_{2}$ 3.85 4.64 74.2 4.00
$p_{3}$ 3.48 3.24 18.7 1.88
Test Set - - 75.1 5.42
Env-2 $p_{2}$ 3.85 4.64 32.5 4.00
$p_{3}$ 3.39 3.20 1.58 1.87
Test Set - - 32.8 5.42
@@ IV Experiments
IV-A Target Distributions Comparison
We compare two target distributions ($p_{2}$, and $p_{3}$) on position error, angular error, collision-with-env rates, and self-collision rates. The results (Tab. I) show $p_{2}$ performs better on all 4 metrics. Especially on collision-with-env and self-collision rates, $p_{3}$ performs badly, even similarly to the test set.
@@ IV-A Target Distributions Comparison
IV-B Runtime Compared with the Classical Method
We compare the runtime of ViIK to the classical method, TRAC-IK+CC. The tolerance of TRAC-IK is set to $1\times 10^{-3}$ which is similar to the errors of ViIK shown in Tab. II. Moreover, Mesh Collider is used for collision checking in TRAC+CC. We let TRAC-IK repeatedly execute until the number of output solutions is the same as ViIK, $t_{ik}$ denotes the total runtime of TRAC-IK. Then we run collision checking for every configuration output by TRAC-IK, $t_{cc}$ denotes the total runtime of collision checking. The runtime of TRAC-IK+CC is:
@@ IV-B Runtime Compared with the Classical Method
TABLE II: Position and Angular Errors on the test set (150,000 random IK problems of Panda manipulator).
@@
IV-C Position Error and Angular Error Compared with the Learning-based Methods
We evaluate ViIK on the test set with 150,000 random IK problems of the Panda manipulator against IKFlow and PaddingFlow. We let ViIK, IKFlow, and PaddingFlow output 1000 solutions for each target pose. In Tab. II, three models (ViIK-2, ViIK-5, and ViIK-10 trained on the data of 2, 5, and 10 scenes respectively) can output competitive solutions compared with existing learning-based methods. ViIK outperforms IKFlow and PaddingFlow on position error which is about 1/2 of IKFlow, but on angular error, ViIK performs worse than PaddingFlow.
TABLE III: Self-collision and Collision-with-Env Rates on the test set (10,000 random IK problems of Panda manipulator).
@@ IV-C Position Error and Angular Error Compared with the Learning-based Methods
IV-D Self-collision and Collision-with-Env Rates
We present ViIK rates of self-collision and collision with the environments on the test sets. We test self-collision and collision-with-env rates on the test set with 10,000 random IK problems of the Panda manipulator. For each target pose, ViIK outputs 1000 solutions. For different scenes, we evaluate ViIK on the same test set but with different images of each scene. We report self-collision and collision-with-env rates over 10,000 target poses, as shown in Tab. III. The self-collision rates of ViIK are lower than 2%, which is significantly lower than the test set. As for the collision-with-env rates, ViIK can perform well in most scenes but is not ideal in some hard scenes (Env-1, and Env-5).
@@ V-B Experimental details
All models are GLOW-based[18] with 24 coupling layers. The training set has 5 million random samples in the C-space with the corresponding end-effector poses and 4882 sets of multi-view images of each environment. The batch size is 1024. The Adamw optimizer is used for training all models. The hyperparameters for training ViIK are shown in Tab. IV. Moreover, the multi-views used in this paper are presented in Fig. 6 with an example in Env-1.
TABLE IV: Hyperparameters for training ViIK.